OpenStudy (anonymous):

Please help❗️❗️😊 Solve for y: 7y-5/3=9y/4

OpenStudy (mathstudent55):

You wrote \(7y - \dfrac{5}{3} = \dfrac{9y}{4} \) Is that your problem or is it the one below? \(\dfrac{7y - 5}{3} = \dfrac{9y}{4}\)

OpenStudy (anonymous):

It is the 2nd one you wrote

OpenStudy (mathstudent55):

Then treat it as proportion and cross multiply it.